# iWalker: Imperative Visual Planning for Walking Humanoid Robot

### v3

@@ V Conclusion & Limitations
We introduced iWalker, a vision-to-control pipeline developed for humanoid robot walking. By incorporating two MPC-based bilevel optimizations through imperative learning, iWalker enables efficient and robust self-supervised training on visual input data, predicting dynamically optimal, collision-free paths and physically feasible footsteps for low-level actuation. Out experiments, conducted in simulation and on real robots, demonstrate the capability of iWalker to navigate through various simulated scenes and unseen real-world environments. However, the system still faces several limitations. The instability of low-level controllers and the limited scale of our platform restrict testing on more challenging terrains, limiting the comprehensiveness of our evaluations. Addressing these issues is critical for bringing our method to more challenging environments.

### v4

@@ V Conclusion & Limitations
We introduced iWalker, a vision-to-control pipeline developed for humanoid robot walking. By incorporating two MPC-based bilevel optimizations through imperative learning, iWalker enables efficient and robust self-supervised training on visual data, predicting dynamically optimal and collision-free paths and physically feasible footsteps for low-level control. Our experiments, conducted in simulation and on real robots, demonstrate the capability of iWalker to navigate through various simulated and real-world environments.
The future work will be on addressing the limitations of the current system design. Particularly, the low-level optimization-based whole-body controllers can yield infeasibility under extreme maneuvers, perhaps due to the limited capability of our small robot platform. This restricts our evaluation of the current work on more challenging environments. Thus, in the next, we will focus on integrating the feasibility of the low-level control, to further enhance the performance of the proposed vision-to-control framework.
Acknowledgements
This work was partially funded by the U.S. Defense Advanced Research Projects Agency (DARPA) grant DARPA-PS-23-13. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of DARPA.